A 15FT LADDER LEANS AGAINST A BULIDING. THE BOTTOM OF THE LADDER IS 7FT FROM THE BUILDING.HOW HIGH IS THE TOP OF THE LADDER?ROUND TO 2 DECIMAL PLACES.
Found 3 solutions by stanbon, asha, jai_kos:
Answer by stanbon(75887)

A 15FT LADDER LEANS AGAINST A BULIDING. THE BOTTOM OF THE LADDER IS 7FT FROM THE BUILDING.HOW HIGH IS THE TOP OF THE LADDER?ROUND TO 2 DECIMAL PLACES.
------------------
Draw the picture.
You have a right triangle.
The hypotenuse is 15 ft.
The base is 7 ft.
EQUATION:
Use Pythagoras to get:
15^2=7^2 + x^2
x^2=176
x=13.27 ft
Cheers,
Stan H.

Answer by asha(30)

length of the ladder is 15 ft
distance from the wall is7ft.
letthe height of the ladder from the bottom be h ft.
using pythagoras theorem we get the eqn.
(15)^2 =7^2 +h^2
225=49 +h^2
h^2=225-49
=176
h= square root of 176
=13.27
the top of the ladder is 13.27 ft. from the ground
good luck!!!

Answer by jai_kos(139)

A
|\
| \
| \
|____\
B C
Let AB be the building, AC be the ladder and
BC is the distance between feet of ladder to the foot of the building.
Given the height of the ladder = 15ft
The distance between feet of ladder to the foot of the building = 7ft.
Using Pythagoras theorem, we find the length of the building which is AB.
AC^2 = AB^2 + BC^2
15 ^2 = AB^2 + 7^2
225 = AB^2 + 49
225 -49 = AB^2
176 = AB^2
Taking the square root, we get
13.26 = AB
Therefore the length of the building is 13.26 feet.